2. Find the general solution of the Eulers equation ty - 3ty + 3y = 0 3. Find the general solution of the Eulers equation
Answer #1

San © Given, {-344 + 3y = 0 t² day 37 dy +39-0 at² at => [t?D2- 3+0+3)9 = 0 (1) Put, t=eZ > z= logit = > dz Differentiating22 mo 3; - 1 So the solution is given by, y= с e3? + C₂ ez ezt =) ēZ=! t and e32= t3 g= att Cz Ane CScanned with CamScanner